OpenStudy (anonymous):

what is the equation in standard form of a parallel line the passes through (0,-2) use y-3=3(x+1) to get answer

OpenStudy (anonymous):

parallel line has the same slope get the y-3=3(x+1) in to y = mx+b m is the slope then use this y-y1 = m(x-x1) and plug in the points (0,-2) , (x1,y2)

OpenStudy (anonymous):

so y+2 = 3(x-0) y = 3x-2 ---> slope intercept form -3x+y = -2----> standard form

OpenStudy (anonymous):

are you still stuck on this?

OpenStudy (anonymous):

yes im trying to see how you got all the answer I have to show my work O///O

OpenStudy (anonymous):